# Safe Text-to-Image Generation: Simply Sanitize the Prompt Embedding

### v1

@@ 4.4 Training Details
Initial Naive Implementation. In our initial naive implementation, we adopt a simple and intuitive objective function aimed at guiding ES to learn the sanitization functionality by minimizing the difference between the sanitized toxic embedding and their corresponding clean embedding. Given a toxic prompt $P_{t}$ and its corresponding clean prompt $P_{c}$, ES’s initial training objective is to transform the sanitized embedding of $P_{t}$ into an approximate embedding of $P_{c}$. We define the sanitized embedding $\hat{\text{Emb}_{t}}$ of $P_{t}$ and the clean embedding $\text{Emb}_{c}$ of $P_{c}$ using the following equations:
$\begin{cases}\hat{\text{Emb}_{t}}=\mathcal{F}_{t}(P_{t})-\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})),\\ \text{Emb}_{c}=\mathcal{F}_{t}(P_{c}),\end{cases}$ (6)
@@
where $\mathcal{F}_{t}$ denotes the text encoder, and $\mathcal{F}_{\theta}$ and $\mathcal{F}_{\phi}$ represent E-Net and S-Net, respectively. Thus, the initial objective function of ES is to minimize the following loss:
$\mathcal{L}_{init}=\min_{\theta,\phi}\|\hat{\text{Emb}_{t}}-\text{Emb}_{c}\|_{2}^{2},$ (7)
@@
where $\mathcal{L}_{init}$ is the mean squared error (MSE) loss, which serves as the objective function penalizing the distance between $\hat{\text{Emb}_{t}}$ and $\text{Emb}_{c}$. Based on the construction of the synthetic data, this objective function encourages ES to approximate target concepts in the embedding toward the anchor concepts while preserving the original predictions of non-target concepts. This design effectively erases potential inappropriate content in the embedding while retaining the integrity of non-target concepts.
@@
However, while the initial objective function Equation 7 provides ES with a learning direction to shift target concepts toward anchor concepts, equipping it with basic sanitization capability, challenges remain in generalizing to synonyms of target concepts. The core issue lies in the difficulty of exhaustively capturing all tokens related to a single concept when constructing synthetic data. For instance, when erasing a pornographic concept, we only define one relevant concept pair, such as $\langle$ nude, dressed $\rangle$. In other words, if a synonym of nude (such as ”bare” as shown in Figure 3) appears in the randomly generated toxic prompt $P_{t}$, the data synthesis strategy would retain it in the clean prompt $P_{c}$ rather than replace it with dressed, as it is not listed in the target concepts. This means that $\mathcal{L}_{init}$ fails to supervise the synonyms of target concepts, subsequently limiting ES’s generalization capability.
@@
Final Improved Implementation. To address this challenge, we design an improved implementation for ES aimed at enhancing the generalization capability to synonyms of target concepts. In the initial implementation, the clean prompt $P_{c}$ consists mainly of random tokens and anchor tokens. We naively use the embedding of $P_{c}$ directly as the supervision target, overlooking the possibility that random tokens in $P_{c}$ may include synonyms of the target concepts. Therefore, considering that $P_{c}$ may still contain inappropriate concepts during its construction, our improved implementation not only sanitizes $P_{t}$ but also sanitizes $P_{c}$ to ensure that $P_{c}$ serves as a cleaner embedding for more accurately supervising ES’s learning. Similar to Equation 6, we define the sanitized embedding of $P_{t}$, denoted as $\hat{\text{Emb}_{t}}$, and the cleaner embedding of $P_{c}$, denoted as $\hat{\text{Emb}_{c}}$, using the following equations:
$\begin{cases}\hat{\text{Emb}_{t}}=\mathcal{F}_{t}(P_{t})-\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})),\\ \hat{\text{Emb}_{c}}=\mathcal{F}_{t}(P_{c})-\beta\cdot\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{c})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{c})),\end{cases}$ (8)
@@
where $\beta$ is a hyperparameter that determines the extent of sanitization applied to $P_{c}$, constrained within the range of [0,1]. The improvement in ES’s generalization to synonymous concepts through the sanitization of $P_{c}$ is primarily driven by the semantic similarity between target concepts and their synonyms. This result arises from the higher cosine similarity between them in the embedding space, where their proximity is relatively small, causing synonymous concepts to typically be located among the neighbors around the target concept [12, 48]. By sanitizing $P_{c}$, this similarity is utilized to identify potential synonymous concept embeddings near the target concept’s embedding. Consequently, potential synonyms in $P_{c}$ can be effectively recognized and transformed into anchor concepts, providing ES with cleaner and more accurate supervision, thereby enhancing its generalization capability for synonymous concepts.
@@
Figure 3 illustrates the final training process of ES. The text encoder of the T2I model first encodes the toxic prompt $P_{t}$ and the clean prompt $P_{c}$ into embedding representations, which are then fed into ES for sanitization. During training, ES shares the same parameters for sanitizing both $P_{t}$ and $P_{c}$. The goal of sanitizing the embedding of $P_{t}$ is to convert its target concepts into anchor concepts. In contrast, the goal of sanitizing the embedding of $P_{c}$ is to convert potential synonyms of target concepts into anchor concepts, thereby providing a cleaner supervision target. Based on the embedding representations in Equation 8, the final objective function is customized as follows:
$\mathcal{L}_{final}=\min_{\theta,\phi}\|\hat{\text{Emb}_{t}}-\hat{\text{Emb}_{c}}\|_{2}^{2},$ (9)
@@
where $\mathcal{L}_{final}$ is the MSE loss that penalizes the distance between $\hat{\text{Emb}_{t}}$ and $\hat{\text{Emb}_{c}}$. This objective function encourages ES to bring potential target concepts and their synonyms in the embedding closer to anchor concepts, further enhancing ES’s robustness.
@@ 5 Evaluation and Analysis
In this section, we begin by outlining the evaluation setting, which includes datasets, safeguard baselines, and evaluation metrics. We then conduct extensive experiments to evaluate the performance of ES across four core dimensions: effectiveness, robustness, specificity, and fidelity. Finally, we explore the impact of hyperparameters on ES’s performance to demonstrate its controllability.
5.1 Evaluation Setup
Dataset. We consider five datasets for evaluating ES. These include a handcrafted prompt dataset, I2P [35], for effectiveness evaluation. Three adversarial prompt datasets, constructed from SneakyPrompt (SP) [49], Ring-A-Bell (RAB) [43], and MMA-Diffusion (MMA) [46], are used for robustness evaluation. Additionally, an image-text pair dataset, COCO-2017[18], is utilized for evaluate specificity and fidelity. We provide a description of the dataset in Appendix A.
Baselines. Considering that ES is a safe generation method, we prioritize comparisons with similar approaches. Specifically, we evaluate six safe generation baselines: NP[1], SLD[35], ESD[7], CA[15], Safe-CLIP[27], and SafeGen[17], along with three additional safeguard baselines: SD-v2.1[41], LG[19], and SC [22]. Unless otherwise specified, we follow prior research [17, 7, 35] by using Stable Diffusion (version 1.4) as the base model for all experiments. We provide a description of the configuration of the baseline method in Appendix B.
Training Setup. We default to using ES to erase the seven inappropriate concepts in I2P unless stated otherwise. The $\langle$ target concept, anchor concept $\rangle$ pairs used in the experiments are as follows:
$\langle$ nude, dressed $\rangle$, $\langle$ illegal, legal $\rangle$, $\langle$ violent, peaceful $\rangle$,
$\langle$ harmful, helpful $\rangle$, $\langle$ shocking, soothing $\rangle$,
where we use antonyms of the seven inappropriate concepts as their anchor concepts. In training ES, we automatically construct training samples using a synthesis strategy based on the above (target concept, anchor concept) pairs, with a batch size of 32 and an AdamW optimizer with a learning rate of $5\times 10^{-4}$ for 10,000 optimization steps. In our experimental setup, the complete training process takes approximately 1 hour of GPU time, thanks to the lightweight architecture design shown in Table II, where ES has only 43.12M parameters, enabling efficient training and inference. Additionally, we set the hyperparameters of ES, $\alpha$ and $\beta$, to 3.0 and 0.9 by default to balance defense performance with generation quality.
Performance Metrics. To comprehensively evaluate ES ’s performance, we analyze it across four dimensions: effectiveness, robustness, specificity, and fidelity, introducing specific metrics for each:
  • Erasure Rate (ER): We use ER as a metric for effectiveness and robustness. ER is calculated by employing an external moderator to detect the presence of inappropriate content in generated images. ER is defined as the reduction of inappropriate concepts in generated images after applying a safeguard. Let $N$ denote the number of inappropriate concepts generated by the original T2I model without any safeguard, and $\hat{N}$ the number after applying the safeguard; then ER is given by $ER=1-\frac{\hat{N}}{N}$. With a higher ER indicating more effective erasure of inappropriate concepts. In line with SafeGEN and SLD, we use the image content detectors to identify inappropriate content, NudeNet[23] for the detection of explicit content and Q16[36] for the detection of other inappropriate content such as violence and hate material.
  • CLIP Score: We use the CLIP Score to evaluate specificity and it calculates cosine similarity between the given text embedding and the image embedding using the CLIP model[29]. For a benign prompt, a higher CLIP Score indicates that the generated image better matches the prompt description, meaning that the safeguard has a minimal effect on unrelated concepts.
  • Frechet Inception Distance (FID) score: We use FID score as a metric to assess fidelity and it is a common metric for evaluating the quality of generated images by measuring the distance between a set of generated images and a reference set. A lower FID score indicates higher semantic similarity between the generated and reference images, thus higher quality. We use the 5,000 real images from the validation set of COCO-2017[18] as the reference set.
Experimental Environment. We implement ES using Python 3.12.4 and Pytorch 2.3.1. All experiments can be conducted on servers running Ubuntu 24.04 LTS, equipped with an Nvidia GeForce RTX 4090 GPUs with 24 GB of VRAM.
Table III: Defensive performance of ES compared to nine baseline methods in terms of the erasure rate of inappropriate content generated from handcrafted crafted and adversarial prompts.
Dataset (Number of Inappropriate Contents) ER (%) $\uparrow$
SD-v2.1 LG SC NP SLD ESD CA Safe-CLIP SafeGen Ours
I2P (Handcrafted Prompt) Hate (91) -2.20 40.66 - 60.44 52.75 54.95 -2.20 49.45 - 70.33
Harassment (266) -5.26 35.34 - 59.40 48.50 42.86 6.77 33.83 - 75.94
Violence (319) 11.29 35.42 - 52.66 49.84 28.84 9.09 43.57 - 84.01
Self-harm (309) 2.91 33.66 - 58.90 59.87 42.39 6.15 46.60 - 84.14
Sexual (674) 33.98 34.27 65.28 48.96 45.25 72.70 16.02 33.23 43.76 93.03
Shocking (423) 2.84 39.72 - 51.06 46.10 40.90 6.38 43.50 - 85.11
Illegal activity (255) 7.06 38.82 - 64.31 52.55 46.67 6.27 45.88 - 80.78
Overall 7.23 36.84 - 56.53 50.69 47.04 6.93 42.29 - 81.91
Adversarial Prompt SP (581) 25.13 39.07 64.37 56.11 49.05 70.57 16.52 38.38 44.23 91.22
RAB (3596) 30.34 49.78 96.77 10.17 3.89 31.89 2.50 42.88 81.23 98.92
MMA (1738) 70.25 77.67 65.88 25.43 9.78 76.01 16.46 79.86 94.53 94.99
Overall 41.91 55.51 75.67 30.57 20.91 59.49 11.83 53.71 73.33 95.04
Figure 4: Comparison of the erasure rates for different exposed body parts between ES and safeguard baselines. SD-v1.4 generated a total of 3596 exposed body parts on the RAB dataset, while our method reduced this number to 39.
5.2 Effectiveness and Robustness Evaluation
We compare ES with nine baseline protection measures, including six safe generation baselines: NP, SLD, ESD, CA, Safe-CLIP, and SafeGen, as well as three additional protection baselines: SD-v2.1, LG, and SC. We evaluate the effectiveness of ES and the baselines in suppressing seven inappropriate concepts on the I2P dataset and evaluate their robustness against adversarial attacks using three adversarial prompt datasets: SP, RAB, and MMA. It is notable that I2P provides parameters such as seed and guidance scale for each prompt to ensure reproducibility of generated content. For fair comparison, we use the additional information provided by the I2P dataset to control image generation. For other prompt datasets, we set a different seed for each prompt to control the generated content, thus reducing errors caused by randomness and fairly reporting the erasure performance of each protective measure under the same random conditions.

### v2

@@ 4.5 Objective Function Formulation
Naive Sanitization Loss. In our initial naive implementation, we adopt a simple and intuitive objective function aimed at guiding ES to learn the sanitization functionality by minimizing the difference between the sanitized toxic embedding and their corresponding anchor embedding. Given a toxic prompt $P_{t}$ and its corresponding anchor prompt $P_{a}$, ES’s naive training objective is to transform the sanitized embedding of $P_{t}$ into an approximate embedding of $P_{a}$. We define the sanitized embedding $\hat{\text{Emb}_{t}}$ of $P_{t}$ and the anchor embedding $\text{Emb}_{a}$ of $P_{a}$ using the following equations:
$\begin{cases}\hat{\text{Emb}_{t}}=\mathcal{F}_{t}(P_{t})-\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})),\\ \text{Emb}_{a}=\mathcal{F}_{t}(P_{a}),\end{cases}$ (5)
@@
where $\mathcal{F}_{t}$ denotes the text encoder, and $\mathcal{F}_{\theta}$ and $\mathcal{F}_{\phi}$ represent E-Net and S-Net, respectively. Thus, the naive objective function of ES is to minimize the following loss:
$\mathcal{L}_{naive}=\min_{\theta,\phi}\|\hat{\text{Emb}_{t}}-\text{Emb}_{a}\|_{2}^{2},$ (6)
@@
where $\mathcal{L}_{naive}$ is the MSE loss, which serves as the objective function penalizing the distance between $\hat{\text{Emb}_{t}}$ and $\text{Emb}_{a}$. This objective function encourages ES to approximate target tokens in the embedding towards the anchor tokens while preserving the original predictions of non-target tokens. This design effectively erases potential inappropriate content in the embedding while retaining the integrity of non-target tokens.
@@
However, while the naive objective function Equation 6 provides ES with a learning direction to shift target tokens toward anchor tokens, equipping it with basic sanitization capability, challenges remain in generalizing to synonyms of target tokens. The core issue lies in the difficulty of exhaustively setting up all the tokens associated with an inappropriate category when constructing synthetic data. For instance, when erasing the pornographic concept, we only define one relevant token pair, such as $\langle$ nude, dressed $\rangle$. In other words, if a synonym of ”nude” (such as ”topless” as shown in Figure 3) appears in the randomly generated toxic prompt $P_{t}$, the data synthesis strategy would retain it in the anchor prompt $P_{a}$ rather than replace it with dressed, as it is not listed in the pre-defined target tokens. This means that $\mathcal{L}_{naive}$ fails to supervise the synonyms of target tokens, subsequently limiting ES’s generalization capability.
@@
Improved Sanitization Loss. To address this challenge, we design an improved implementation for ES aimed at enhancing the generalization capability to synonyms of target tokens. Considering that $P_{a}$ may still contain inappropriate tokens during its construction, our improved implementation not only sanitizes $P_{t}$ but also sanitizes $P_{a}$ to ensure that $P_{a}$ serves as a cleaner embedding for more accurately supervising ES’s learning. Similar to Equation 5, we define the sanitized embedding of $P_{t}$, denoted as $\hat{\text{Emb}_{t}}$, and the cleaner anchor embedding of $P_{a}$, denoted as $\hat{\text{Emb}_{a}}$, using the following equations:
$\begin{cases}\hat{\text{Emb}_{t}}=\mathcal{F}_{t}(P_{t})-\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})),\\ \hat{\text{Emb}_{a}}=\mathcal{F}_{t}(P_{a})-\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{a})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{a})),\end{cases}$ (7)
@@
The improvement in ES’s generalization to synonymous tokens through the sanitization of $P_{a}$ is primarily driven by the semantic similarity between target tokens and their synonyms. This result arises from the higher cosine similarity between them in the embedding space, where their proximity is relatively small, causing synonymous tokens to typically be located among neighbors around the target token [16, 52]. By sanitizing $P_{a}$, this similarity is utilized to identify potential synonymous token embeddings near the target token’s embedding. Consequently, potential synonyms in $P_{a}$ can be effectively recognized and transformed into anchor tokens, providing ES with cleaner and more accurate supervision, thereby enhancing its coverage of inappropriate concepts.
@@
Figure 3 illustrates the training process for the improved implementation. The text encoder of the T2I model first encodes the toxic prompt $P_{t}$ and the anchor prompt $P_{a}$ into embedding representations, which are then fed into ES for sanitization. During training, ES shares the same weights for sanitizing both $P_{t}$ and $P_{a}$. The goal of sanitizing the embedding of $P_{t}$ is to convert its target tokens into anchor tokens. In contrast, the goal of sanitizing the embedding of $P_{a}$ is to convert potential synonyms of target tokens into anchor tokens, thereby providing a cleaner supervision target. Based on the embedding representations in Equation 7, the final objective function is customized as follows:
$\mathcal{L}_{sanitize}=\min_{\theta,\phi}\|\hat{\text{Emb}_{t}}-\hat{\text{Emb}_{a}}\|_{2}^{2},$ (8)
@@
where $\mathcal{L}_{sanitize}$ is the MSE loss that penalizes the distance between $\hat{\text{Emb}_{t}}$ and $\hat{\text{Emb}_{a}}$. This objective function encourages ES to bring potential target tokens and their synonyms in the embedding closer to anchor tokens, further enhancing the robustness of ES by improving its generalization.
Benign Loss. In summary, $\mathcal{L}_{sanitize}$ primarily guides ES in learning how to sanitize embeddings, but it does not sufficiently address the handling of benign prompts. To balance defense effectiveness and generation quality, we introduce a benign loss term that helps ES preserve the semantic quality of outputs when processing benign prompts. Specifically, we use image captions from the COCO-2017 training set [22] as benign prompts $P_{b}$, and design a corresponding benign loss function to encourage ES to produce sanitized embeddings that remain semantically consistent with the original inputs. This loss is formalized as:
$\mathcal{L}_{benign}=\min_{\theta,\phi}\|\hat{\text{Emb}_{b}}-\text{Emb}_{b}\|_{2}^{2},$ (9)
where $\text{Emb}_{b}=\mathcal{F}_{t}(P_{b})$ denotes the original embedding of the benign prompt, and $\hat{\text{Emb}_{b}}$ is the sanitized embedding produced by ES. Accordingly, the final objective function of ES is defined as:
$\mathcal{L}_{final}=\mathcal{L}_{sanitize}+\lambda\cdot\mathcal{L}_{benign},$ (10)
where $\lambda$ is a balancing coefficient between the two loss terms. Our ablation study demonstrates that setting $\lambda=0.01$ achieves an effective trade-off between defense capability and generation quality.
5 Evaluation Setup
Dataset. We use five datasets to comprehensively evaluate ES. The I2P dataset [38], a manually curated collection of prompts covering seven categories of inappropriate content, serves to assess the effectiveness of the defense. Additionally, we adopt three adversarial prompt datasets generated by attack methods — SneakyPrompt (SP) [52], Ring-A-Bell (RAB) [47], and MMA-Diffusion (MMA) [50] — to evaluate the robustness of ES. Finally, we use the benign image-text pair from the COCO-2017 [22] validation set to assess the specificity and fidelity of ES. Detailed descriptions of these datasets are provided in Appendix A.
Baselines. Considering that ES is a safe generation method, we prioritize comparisons with similar approaches. Specifically, we evaluate eight safe generation baselines: NP [1], SLD [38], ESD [8], POSI [49], Safe-CLIP [30], SafeGen [21], Rickrolling (RR) [46], and Moderator [48], along with three additional safeguard baselines: SD-v2.1 [44], LG [23], and SC [26]. Unless otherwise specified, we follow previous research [21, 8, 38] using Stable Diffusion (version 1.4) as the base model for all experiments. We provide a description of the configuration of the baseline methods in Appendix B.
Training Setup. I2P defines inappropriate content based on the findings of Gebru et al. [9], categorizing it into seven types: hate, harassment, violence, self-harm, sexual content, shocking images, and illegal activities. Following I2P’s definitions of these seven types of inappropriate content, we use ES to erase these concepts and define a set of $\langle$ target token, anchor token $\rangle$ pairs for each category:
$\langle$ violent, peaceful $\rangle$, $\langle$ harmful, helpful $\rangle$, $\langle$ nude, dressed $\rangle$,
$\langle$ shocking, soothing $\rangle$, $\langle$ illegal, legal $\rangle$.
In all subsequent experiments, we use a token pair for each inappropriate category to eliminate all seven types of inappropriate content simultaneously. The selection of token pairs can be flexibly adjusted on the basis of practical requirements, and increasing the number of token pairs holds the promise of further improving the diversity of synthesized data. Since we fixed the amount of synthetic data used for training, adjusting the number of token pairs does not affect the training overhead.
To train ES, we automatically construct training samples using a synthesis strategy based on the above seven token pairs, with a batch size of 32 and an AdamW optimizer with a learning rate of $5\times 10^{-4}$ for 10,000 optimization steps, resulting in a total of 0.32M training samples. In our experimental setup, the complete training process takes approximately 1 hour of GPU time, thanks to the lightweight architecture design shown in Table II, where ES has only 43.12M parameters, allowing efficient training and inference. Additionally, we set the hyperparameter $\alpha$ to 3.0 by default to balance defense performance and generation quality.
Performance Metrics. To evaluate ES ’s performance, we analyze it across four dimensions: effectiveness, robustness, specificity, and fidelity, with corresponding metrics for each:
  • Erasure Rate (ER): We use ER as a metric for effectiveness and robustness. ER is defined as the reduction of inappropriate concepts in generated images after applying a safeguard. Let $N$ denote the number of inappropriate concepts generated by the original T2I model without any safeguard, and $\hat{N}$ the number after applying the safeguard; then ER is given by $ER=1-\frac{\hat{N}}{N}$. With a higher ER indicating a more effective erasure of inappropriate concepts. Consistent with prior studies [21, 38, 8], we employ image content detectors to identify inappropriate content. Specifically, we use NudeNet[27] with the score threshold set to 0 to detect explicit content, and Q16[39] to identify other types of inappropriate content, such as violence and hate.
  • CLIP Score: We use the CLIP score to evaluate specificity, which is calculated by the CLIP model [32] based on the cosine similarity between pairs of benign text and image embeddings. A higher CLIP score indicates greater consistency between the generated image and the prompt description, suggesting that the safeguard has a smaller impact on the semantic consistency of benign concept generation. We calculate the CLIP score using benign prompts from the COCO-2017 validation set paired with generated images.
  • Frechet Inception Distance (FID) Score: We use the FID score to evaluate fidelity, which is a common metric for assessing the quality of generated images by measuring the distance between a set of generated images and a reference set. A lower FID score indicates higher semantic similarity between the generated images and the reference images, thus indicating higher quality. We calculate the FID score using 5,000 real images from the COCO-2017 validation set and 5,000 generated images based on benign prompts from this validation set.
Experimental Environment. We implement ES using Python 3.12.4 and Pytorch 2.3.1. All experiments are conducted on an Ubuntu 24.04 LTS server equipped with a Nvidia GeForce RTX 4090 GPU and 24 GB of VRAM.
Table III: [RQ1] Defensive performance of ES compared to eleven baseline methods in terms of the erasure rate of inappropriate content generated from handcrafted crafted and adversarial prompts. The red values indicate that more inappropriate content is generated compared to the original SD-v1.4 model. The bolded values indicate the highest performance. The underlined italicized values are the second highest performance.
Dataset (Number of Inappropriate Contents) ER (%) $\uparrow$
SD-v2.1 LG SC NP SLD ESD POSI Safe-CLIP SafeGen RR Moderator Ours
I2P (Handcrafted Prompt) Hate (91) -2.20 40.66 - 60.44 52.75 54.95 41.78 49.45 - -3.29 - 68.13
Harassment (266) -5.26 35.34 - 59.40 48.50 42.86 28.46 33.83 - -1.87 - 67.45
Violence (319) 11.29 35.42 - 52.66 49.84 28.84 34.79 43.57 - 0.00 - 81.19
Self-harm (309) 2.91 33.66 - 58.90 59.87 42.39 33.87 46.60 - 1.61 - 70.65
Sexual (674) 33.98 34.27 65.28 48.96 45.25 72.70 22.55 33.23 43.76 4.01 50.29 90.21
Shocking (423) 2.84 39.72 - 51.06 46.10 40.90 35.84 43.50 - 2.35 - 79.25
Illegal activity (255) 7.06 38.82 - 64.31 52.55 46.67 41.73 45.88 - -2.36 - 66.54
Overall 7.23 36.84 - 56.53 50.69 47.04 34.15 42.29 - 0.06 - 74.77
Adversarial Prompt SP (581) 25.13 39.07 64.37 56.11 49.05 70.57 27.37 38.38 44.23 -5.51 55.93 90.88
RAB (3596) 30.34 49.78 96.77 10.17 3.89 31.89 25.12 42.88 81.23 1.83 40.48 99.14
MMA (1738) 70.25 77.67 65.88 25.43 9.78 76.01 66.11 79.86 94.53 -2.18 26.29 85.16
Overall 41.91 55.51 75.67 30.57 20.91 59.49 39.53 53.71 73.33 -1.95 40.90 91.73
@@ 6 Evaluation and Analysis
We design and conduct a series of comprehensive experiments to systematically answer the following research questions (RQ).
  • $[RQ1]$ How effective is ES in mitigating unsafe content generation caused by manual prompts or adversarial prompts?
  • $[RQ2]$ How well does ES preserve the quality of benign content generation?
  • $[RQ3]$ Why is ES able to identify inappropriate tokens that are not explicitly annotated during training?
  • $[RQ4]$ How does ES perform when combined with other safeguards or different diffusion models?
6.1 RQ1: Defense Performance Evaluation
We compare ES with eleven baseline safeguards, including eight safe generation baselines: NP, SLD, ESD, POSI, Safe-CLIP, SafeGen, RR, and Moderator, as well as three additional protection baselines: SD-v2.1, LG, and SC. It is important to note that SC and SafeGen are only applicable to explicit content, and we evaluate them solely within this category. Furthermore, considering that Moderator experiences significant performance degradation when erasing multiple concepts, we use it exclusively for handling explicit content and evaluate it accordingly. We assess the effectiveness of ES and its baselines in suppressing seven types of inappropriate concepts on the I2P dataset. Additionally, we evaluate their robustness against adversarial attacks using three adversarial prompt datasets: SP, RAB, and MMA. To minimize the impact of randomness, we assign a fixed random seed to each prompt.